# High-performance superconducting two-qubit gate using a three-mode tunable coupler

### v1

@@ I Introduction
Development of superconducting qubit-based quantum computing devices has achieved a number of major milestones during the past few years [1, 2, 3, 4, 5]. The common theme is the usage of transmon qubits [6] with dispersive readout and capacitive coupling [7]. Nevertheless, the relatively large two-qubit gate error rates undermine both the practical application of such machines as well as the proof-of-principle advantage demonstration [8].
@@
For implementation of two-qubit gates, there has been a variety of different approaches, each with its strengths and weaknesses. The simplest scheme involves resonant exchange of qubit populations triggered by a frequency shift of one of the qubits [9, 10]. For transmons, one can choose between an iSWAP-type gate [11, 12, 13] or a CPHASE-type gate [12, 13, 14]. The gate duration can be short and is determined by the qubit-qubit coupling constant. This scheme is prone to (i) significant qubit dephasing, because the qubits are operated outside of their sweet spots, which can be addressed with net-zero pulses [15], (ii) crossing of two-level systems (TLS) defects during gate operation [16], and (iii) large residual ZZ crosstalk when the qubits are detuned from each other. The need to traverse a large frequency range with the qubit and thus potentially cross strongly coupled TLS defects can be circumvented, for example, using a parametrically modulated signal [17]. For parametrically activated gates, however, the effective coupling strength is only a fraction of the underlying coupling between qubits. Moreover, the qubit is still detuned from it’s sweet spot, resulting in additional dephasing, which can be reduced by applying a modulation around the DC sweet spot [18].
@@
Another approach is by using a tunable coupler element, either for the parametric modulation gate [19, 20] or direct resonance gate [21, 22]. For CPHASE-type gates one can also use an adiabatic flux pulse, which results in slower gates for the same coupling strength [23, 24, 25, 26, 27], but does not require precise resonance between the qubits. Otherwise, one must tune both the qubit frequency and the coupler. To avoid dephasing, one can use an AC signal to tune the effective qubit frequency for the duration of the gate [28]. By controlling both qubit frequencies and coupler flux with DC signals, one can implement the full continuous gate set that, up to single-qubit gates, includes all excitation-number-preserving unitaries [12]. For tunable couplers, one of the main desired features is cancelling out crosstalk. This can be done in a straightforward manner by using destructive interference between two coupling paths. These coupling paths can either be of Fano-type, with one coupling being resonant [29] or both coupling paths having resonances [30]. In either case, the residual interaction strongly depends on the relation between qubit and coupler frequencies.
In this work, we present a unitary cell suitable for a scalable quantum processor architecture and consisting of a two concentric transmons [31, 32, 17, 33, 34] and a tunable three-mode coupler for a controlled ZZ interaction. The coupler consists of a coplanar shorted by a SQUID in the middle and Josephson junctions (JJs) at the coplanar ends, being an upgrade of a promising tunable coupler [30, 35, 29]. Two JJs at the coplanar ends are implemented in order to suppress a parasitic ZZ interaction. Thus, by tuning the magnetic flux in the coupler SQUID, we are able to vary the interaction strength between qubits by two orders of magnitude. A native two-qubit CZ gate is implemented by leveraging a cross-Kerr interaction [30, 36, 35, 29, 37]. Our experimental CZ cross-entropy benchmarking (XEB) repeatedly gives the gate fidelity above 98%, being limited mostly by the qubits coherence time.
@@ II Device description
The two-qubit device explored in this work is shown in Fig. 1. It consists of two tunable transmons connected by a tunable three-mode coupler with an individual flux-control line. We utilize here concentric weakly flux-tunable transmon qubits [38] for reduction of dielectric losses, containing a 3-JJs circuit [39] for reducing flux noise sensitivity at the expence of narrowing tunability range, shown in Fig. 1. Our 3-JJs qubit circuit consists of two JJs forming an asymmetric DC SQUID connected in series to a third JJ. Each qubit has a single flux-coupled control line for both dc and microwave manipulations. The physical pitch between qubits is 2 mm. Capacitive coupling between the transmon and the coupler is provided by arc-shaped electrodes [17, 40]. The three-mode coupler, that we present in the current work, has three nodes and consists of a coplanar waveguide segment, terminated with JJs, and shorted in the middle by a DC SQUID, see Fig. 1(b).
The Hamiltonian of the whole qubit-coupler-qubit system can be written as:
$\hat{H}_{\text{int}}=\sum_{i\neq j}E_{Cij}\hat{n}_{i}\hat{n}_{j},$ (2)
where $\{i,j\}$ is the set of circuit nodes pairs and $E_{Cij}$ is the energy of the capacitive interaction. Exact expression of the Hamiltonian (1) is given in Appendix A.
@@
Figure 1: The two-qubit cell. (a) An optical image of a sample. Each transmon (red) has a round island grounded through a circuit of three Josephson junctions: one junction in series with a SQUID (yellow). Each qubit has a separate control line both for tuning and excitation. The transmons are connected to each other through a tunable coupler (blue). The coupler shares an arc-shaped mutual capacitor with each qubit. The coupler is realised by a coplanar line with a SQUID in the middle. A separate line controls a magnetic flux through coupler’s SQUID providing tunability. In order to suppress parasitic ZZ interaction, an additional Josephson junction in each end of the coplanar is embedded. (b) Three key elements of the cell (from left to right): the three Josephson junction circuit of the transmon, the Josephson junction at one of the coplanar ends, the SQUID of the coupler. (c) A principal electrical circuit of the two-qubit cell. The two transmons are shown in red, the coupler is shown in blue. The more detailed electrical circuit can be found in Appendix A.
@@
Figure 2: The measurement of the ZZ interaction. (a) The pulse sequence performed on our device: a Ramsey-type experiment on the qubit B while control qubit A is initialized in its ground or excited state, the coupler is tuned by a Hann-shaped flux pulse of the duration $t_{d}$. The argument $\varphi$ of the virtual $R_{Z}$ gate is determined by an offset frequency $f_{\text{offset}}$ and pulse duration $t_{d}$. (b, c) The resulted pattern of the qubit B relative excitation. The red line indicates the boundary between two different offset frequencies used in the experiments. (d) The ZZ interaction strength defined as the difference between the oscillation frequencies of the qubit B population in (b) and (c). The red dashed line shows the ZZ interaction strength $\zeta_{ZZ}$, calculated using the full-circuit model described in Appendix A with the experimental parameters obtained by two-tone spectroscopy measurements of each computational qubit and the coupler. The optimized parameters are given in Appendix A Table 3.
@@ III.1 ZZ interaction strength
During the study of the ZZ interaction between qubits, both transmons are kept in the upper flux sweet spot, while the magnetic flux in the coupler loop is varied between zero flux and half flux quantum $\Phi_{\mathrm{ext}}\in\left[0,\Phi_{0}/2\right]$. In each subsystem, we take into account the first five energy levels. By numerical diagonalization of the system Hamiltonian (1) we analitically calculate a ZZ-interaction strength shown in Fig. 2(d) together with experimental data. Details on the numerical calculation can be found in Appendix A.
@@
In our experiment we measure the strength of the ZZ-interaction between qubits as a function of an external flux through the coupler SQUID. We use a Ramsey-type pulse sequence, consisting of two $\pi/2$ pulses applied to the qubit B, and a Hann-shaped flux pulse of duration $t_{d}$, applied to the coupler, between them, as depicted in Fig. 2(a). We activate [Fig. 2(c)] or do not activate [Fig. 2(b)] the ZZ interaction with an $X$ operation on the control qubit A and measure the population of the target qubit B. For each voltage amplitude generating a flux through coupler SQUID, we find the frequency of population oscillations as a function of the delay $t_{d}$ for both cases when the qubit A remains at the ground state [Fig. 2(b)] and goes to the excited state [Fig. 2(c)]. We define the strength of the ZZ interaction $\zeta_{ZZ}$ as the difference of these frequencies. At this point we should mention, that in order to observe oscillations on scales of hundreds of nanoseconds even at low external flux, we apply virtual rotation $R_{Z}(\varphi)$ [41] about Z axis on the qubit B before the second $R_{X}(\pi/2)$ gate. The phase $\varphi$ depends on the duration of the coupler pulse $\varphi=2\pi f_{\text{offset}}t_{d}$, where $f_{\text{offset}}$ is a defined offset from the frequency of the qubit B. For convenience, we take two offset frequencies equal to 10 MHz for low and 50 MHz for high CZ-pulse amplitudes. The border of the offset frequency change is denoted with the red line in Fig. 2(b, c).
@@
We provide the measured ZZ interaction strength in Fig. 2(d) as blue points. We extract system parameters from the two-tone spectroscopy of both qubits. Knowing the qubit frequencies $f_{\text{A}},f_{\text{B}}$ and anharmonicities $\alpha_{\text{A}},\alpha_{\text{B}}$ as functions of qubit external flux, one can get a critical current of two 3-JJs circuits of the qubits, marked with the red color in Fig. 1. From the two-tone spectroscopy of the coupler we obtain the coupler frequency $f_{\text{C}}$ at external magnetic flux $\Phi_{\mathrm{ext}}=0.5\Phi_{0}$. Using the full system Hamiltonian with obtained qubit JJs parameters and capacitances from the design, we fit the four JJs related to the coupler, colored blue in Fig. 1(c). We optimize the three critical currents of the coupler junctions in order to meet an experimental values of the ZZ interaction at the zero external magnetic flux $\zeta_{ZZ}(\Phi_{\text{ext}}=0)$ and the coupler frequency $f_{\text{C}}$. The designed and fitted parameters required for the system Hamiltonian (1) are presented in Tables 2 and 3 correspondingly. Also, we should mention that all junctions were produced in one fabrication stage, having the same critical current density. Thus, during the optimization, we vary only the JJs areas, assuming the same critical current density for all of them. Finally, the calculated the ZZ interaction strength as a function of the external magnetic flux $\Phi_{\text{ext}}$ plotted in Fig. 2(d).
@@
The residual ZZ interaction at $\Phi_{\text{ext}}=0$, i.e. when the coupler is deactivated, was obtained to be 416 kHz, which is almost 300 kHz higher then expected. The three-mode tunable coupler calculated for the designed parameters gives the ZZ interaction strength $\zeta_{ZZ}$ below 150 kHz (see Fig. 7(b) in Appendix A). Such a difference in $\zeta_{\mathrm{ZZ}}$ is caused by fabrication imperfections, i.e. by the difference between the design and the fabricated device. However, the difference between expected and approximated parameters does not significantly influence the overall gate performance, as we show in further simulations and experiments.
Figure 3: The CZ gate simulation for the experimentally obtained device parameters. (a) The fidelity $F$ (red), the target population $P_{t}$ calculated as a mean projection of the computational states on themselves during the CZ gate (dashed blue) and the high-energy levels leakage error $L_{e}$ (dashed-dotted green) are shown as functions of the CZ gate duration. (b) The corresponding amplitude of the external flux in the coupler’s SQUID. (c) For the gate duration of 43.3 ns (star in panel (a)) the time evolution of the conditional two-qubit phase $\theta$ (red), the target population $P_{t}$ (dashed blue), and the leakage error $L_{e}$ (dashed-dotted green) are simulated. The fidelity estimated by Eq. (3) is $F=0.999(7)$. (d) The coupler flux pulse with Hann-shaped edges is used both in simulations and experiments.
III.2 CZ gate concept

### v2

# Three-mode tunable coupler for superconducting two-qubit gates

@@ I Introduction
Development of superconducting qubit-based quantum computing devices has achieved a number of major milestones during the past few years [1, 2, 3, 4, 5]. The common theme is the usage of transmon qubits [6] with dispersive readout and capacitive coupling [7]. Nevertheless, the relatively large two-qubit gate error rates remain to be the limiting factor for both the practical application of such machines as well as the proof-of-principle advantage demonstrations [8].
@@
For transmon qubits, there is variety of different implementations of two-qubit gates, each with its strengths and weaknesses. The simplest scheme involves resonant exchange of qubit populations triggered by a frequency shift of one of the qubits [9, 10]. For transmons, one can choose between an iSWAP-type gate [11, 12, 13] or a CPHASE-type gate [12, 13, 14]. The gate duration can be short and is determined by the qubit-qubit coupling constant. This scheme is prone to (i) significant qubit dephasing, because the qubits are operated outside of their sweet spots, which can be addressed with net-zero pulses [15], (ii) crossing of two-level system (TLS) defects during gate operation [16], and (iii) large residual ZZ crosstalk when the qubits are detuned from each other. The need to traverse a large frequency range with the qubit and thus potentially cross strongly coupled TLS defects can be circumvented, for example, using a parametrically modulated signal [17]. For parametrically activated gates, however, the effective coupling strength is only a fraction of the underlying coupling between qubits. Moreover, the qubit is still detuned from its sweet spot, resulting in additional dephasing, which can be reduced by applying a modulation around the DC sweet spot [18].
@@
Another approach is based on tunable coupler elements, either for the parametric modulation gate [19, 20] or direct resonance gate [21, 22]. For CPHASE-type gates one can also use an adiabatic flux pulse, which results in slower gates for the same coupling strength [23, 24, 25, 26, 27], but does not require precise resonance between the qubits. Otherwise, one must tune both the qubit frequency and the coupler. To avoid dephasing, one can use an AC signal to tune the effective qubit frequency for the duration of the gate [28]. By controlling both qubit frequencies and coupler flux with DC signals, one can implement the full continuous gate set that, up to single-qubit gates, includes all excitation-number-preserving unitaries [12]. For tunable couplers, one of the main desired features is cancelling out crosstalk. This can be done in a straightforward manner by using destructive interference between two coupling paths. These coupling paths can either be of Fano-type, with one coupling being resonant [29, 30, 31] or both coupling paths having resonances [32]. In either case, the residual interaction strongly depends on the relation between qubit and coupler frequencies.
To address the issue of a relatively large residual ZZ interaction, several techniques can be identified capable of mitigating this unwanted effect. Among the most common methods are passive suppression using highly detuned qubits [23, 33] or possessing opposite-sign anharmonicities [34, 35], as well as active dynamical decoupling protocols [36, 37], which are widely employed in experiments involving long idling times [3, 4]. Another promising strategy involves utilizing the AC Stark shift effect, which has demonstrated encouraging performance in recent experiments [38, 39]. Alternatively, similar suppression can be achieved by applying a weak microwave drive to the coupler [40].
In this work, we present a building block suitable for a scalable quantum processor architecture and consisting of a two concentric transmons [41, 42, 17, 43, 44] and a tunable three-mode coupler for a controlled ZZ interaction. The distinctive feature of the three-mode coupler is its lack of dependence on high asymmetry in the Josephson junctions of the coupler SQUID. This characteristic provides a significant advantage in cases where achieving high reproducibility of junctions with different sizes is challenging. For instance, in the case of the double-transmon coupler type [27], the achieved asymmetry is critical, as deviations can lead to significant changes in the energy structure of the flux qubit. Similarly, a transmon-type coupler [23, 24, 29, 32] requires precise control of asymmetry to simultaneously achieve both a small residual ZZ interaction and a coupling strength on the order of tens of megahertz at the operating point. The proposed three-mode coupler achieves comparable levels of residual ZZ interaction strength while providing a significantly larger tunable range compared to conventional single-mode couplers, assuming identical design parameters. By tuning the magnetic flux in the coupler SQUID, we are able to vary the interaction strength between qubits by two orders of magnitude. We implement a native two-qubit CZ gate by leveraging a cross-Kerr interaction [23, 24, 27, 45, 46] and estimate a gate fidelity above 98% using cross-entropy benchmarking, limited mostly by the qubits’ coherence time.
@@ II Device description
The two-qubit device consists of two tunable transmons connected by a tunable three-mode coupler with an individual flux-control line. We utilize concentric weakly flux-tunable transmon qubits [47] for reduction of dielectric losses, containing a three-Josephson-junction (3-JJ) circuit [48] for reducing flux noise sensitivity at the expence of narrowing tunability range, shown in Fig. 1. Our 3-JJ qubit circuit consists of two Josephson junctions (JJs) forming an asymmetric DC SQUID connected in series to a third JJ. Each qubit has a single flux-coupled control line for both dc and microwave signals. The physical pitch between qubits is 2 mm. Capacitive coupling between the transmon and the coupler is provided by arc-shaped electrodes [17, 49]. The three-mode coupler has three nodes and consists of a coplanar waveguide segment, terminated with JJs, and shorted in the middle by a DC SQUID, see Fig. 1(b). The Hamiltonian of the system and the circuit parameters are described in Appendix A. Due to the mirror-symmetry of the circuit, the eigenmodes are either even or odd (Fig. 1(d)). The even modes are flux-tunable. The odd mode is not flux-tunable, since there is no currents flowing through the SQUID associated with it.
Al/AlOx/Al Josephson junctions in the device are formed using a standard Dolan bridge technique [50] in a single vacuum cycle, ensuring they share the same critical current density. The galvanic contact with the ground plane is achieved by depositing bandages through a single-layer organic mask following argon milling of the aluminum oxide [51]. More details on the fabrication can be found in [52].
@@
Figure 1: The two-qubit cell. (a) Optical image of the sample. Each transmon (red) has a round island grounded through a circuit of three Josephson junctions: a single junction in series with a SQUID (yellow). Each qubit has a separate flux control line both for frequency tuning and excitation. The transmons are connected to each other through a tunable coupler (blue). The coupler is realised by a coplanar line with a SQUID in the middle. At either end of the coupler, there is an arc-shaped electrode with a large mutual capacitance with the respective qubit island. A separate line controls the magnetic flux through the coupler’s SQUID, providing frequency tunability. Additional Josephson junctions are embedded at each end of the coplanar. (b) Three key elements of the cell (from left to right): the three Josephson junction circuit of the transmon, the Josephson junction at one of the coplanar ends, the SQUID of the coupler. (c) Simplified electrical circuit schematic of the two-qubit cell. The two transmons are shown in red, the coupler is shown in blue. The more detailed electrical circuit can be found in Appendix A. (d) Numerically calculated coupler energy levels as a function of the external magnetic flux in the coupler SQUID (see Appendix A). The normal modes of the coupler are determined by the states of its three nodes. Arrows indicate signs of charge oscillations across these nodes for each mode.
@@
Figure 2: The measurement of the ZZ interaction. (a) The pulse sequence performed on our device: a Ramsey-type experiment on the qubit B while control qubit A is initialized in its ground or excited state, the coupler is tuned by a Hann-shaped flux pulse of the duration $t_{d}$. The argument $\varphi$ of the virtual $R_{Z}$ gate is determined by an offset frequency $f_{\text{offset}}$ and pulse duration $t_{d}$. (b, c) The resulted pattern of the qubit B relative excitation. The red line indicates the boundary between two different offset frequencies used in the experiments. (d) The ZZ interaction strength defined as the difference between the oscillation frequencies of the qubit B population in (b) and (c) (blue dots). The red dashed line shows the ZZ interaction strength $\zeta_{ZZ}$, calculated using the full-circuit model described in Appendix A with the experimental parameters obtained by two-tone spectroscopy measurements of each computational qubit and the coupler. The optimized parameters are given in Appendix A Table 2. The blue dashed and green dashed-dotted lines represent the ZZ interaction strengths, $\zeta_{ZZ}$ and the effective XX coupling, $J_{\mathrm{XX}}$, calculated using the approach described in Section III.2 and Appendix B, with the experimental parameters as well. The idle bias point corresponds to the zero external magnetic flux, where the smallest ZZ interaction between qubits is achieved.
@@ III.1 ZZ interaction strength
In our experiment we measure the strength of the ZZ interaction between qubits as a function of an external flux through the coupler SQUID. We use a Ramsey-type pulse sequence, consisting of two $\pi/2$ pulses applied to the qubit B, and a Hann-shaped flux pulse of duration $t_{d}$, applied to the coupler, between them, as depicted in Fig. 2(a). The results of the Ramsey experiment on qubit B as a function of flux pulse amplitude and length are shown in Fig. 2(b) for de-excited qubit A and in Fig. 2(c) for excited qubit A. For each voltage amplitude generating a flux through coupler SQUID, we find the frequency of population oscillations as a function of the delay $t_{d}$. We define the strength of the ZZ interaction $\zeta_{ZZ}$ as the difference of these frequencies. In order to observe oscillations on scales of hundreds of nanoseconds even at low external flux, we apply a virtual rotation $R_{Z}(\varphi)$ [53] about the Z axis on qubit B before the second $R_{X}(\pi/2)$ gate. The phase $\varphi$ depends on the duration of the coupler pulse $\varphi=2\pi f_{\text{offset}}t_{d}$, where $f_{\text{offset}}$ is a defined offset from the frequency of the qubit B. For convenience, we take two offset frequencies equal to 10 MHz for low and 50 MHz for high CZ-pulse amplitudes. The red line in Fig. 2(b, c) shows the boundary between the measurements with the different offset frequencies.
@@
The measured ZZ interaction strength are shown in Fig. 2(d) with blue points. We extract system parameters from the two-tone spectroscopy of both qubits. Knowing the qubit frequencies $f_{\mathrm{q_{A}}},f_{\mathrm{q_{B}}}$ and anharmonicities $\delta_{\mathrm{q_{A}}},\delta_{\mathrm{q_{B}}}$ as functions of qubit external flux, we estimate the critical currents of the six Josephson junctions of the qubit circuits with least-squares fitting. From the two-tone spectroscopy of the coupler we obtain the coupler frequency $f_{\text{C}}$ at external magnetic flux $\Phi_{\mathrm{ext}}=0.5\Phi_{0}$. Using the full system Hamiltonian with obtained qubit JJ critical currents and capacitances from the design, we estimate the critical currents of the four JJs related to the coupler, colored blue in Fig. 1(c). In the detailed scheme in Fig. 5 these junctions are noted as $J_{5},J_{6}$ and $J_{4}=J_{7}$. The last two junctions are considered to be equal. We optimize the three critical currents of the coupler junctions in order to meet an experimental values of the ZZ interaction at the zero external magnetic flux $\zeta_{ZZ}(\Phi_{\text{ext}}=0)$ and the coupler frequency $f_{\text{C}}$. The designed and fitted parameters of the system Hamiltonian (21) are presented in Tab. 2. In the optimization procedure, we vary only the JJs areas, assuming the same critical current density for all of them.
During the study of the ZZ interaction between qubits, both transmons are kept in the upper flux sweet spot, while the magnetic flux in the coupler loop is varied between zero flux and half flux quantum $\Phi_{\mathrm{ext}}\in\left[0,\Phi_{0}/2\right]$. In each subsystem, we take into account the first five energy levels. The ZZ interaction strength, $\zeta_{ZZ}$, is calculated by the system Hamiltonian (21) numerical diagonalization and is defined as the frequency difference between the state in which both computational qubits are excited $|ee\rangle$ and the sum of the frequencies corresponding to the states where only one qubit is excited, $|eg\rangle$ and $|ge\rangle$: $\zeta_{ZZ}=f_{ee}-f_{eg}-f_{ge}$, assuming the ground state frequency is set to zero. The simulated dependence of $\zeta_{ZZ}$ on the external magnetic flux $\Phi_{\text{ext}}$ is shown in Fig. 2(d), together with the experimental data. We denote the states of the computational qubits as $|g\rangle,|e\rangle,|f\rangle$ and use numerical notation for coupler modes. For simplicity, we omit explicit indices for the coupler state and assume projection onto its ground state, for instance, the state $|ee\rangle$ corresponds to $|ee000\rangle$. Details on the numerical calculation can be found in Appendix A.
@@
At the idle bias point of the coupler ($\Phi_{\mathrm{ext}}=0$), we obtain a residual ZZ interaction constant of 416 kHz. This by far exceeds the target value of 146 kHz obtained from calculations with the design parameters (see Fig. 6(b) in Appendix A). Such a difference in $\zeta_{\mathrm{ZZ}}$ is caused by fabrication imperfections, i.e. by the difference between the design and the fabricated device. However, the difference between expected and approximated parameters does not significantly influence the overall gate performance, as we show in further simulations and experiments.
III.2 Linear-response theory for qubit-qubit coupling
In the linear response regime, it is possible to derive an analytical expression for the XX interaction strength between the qubits. Given the capacitance matrix and the diagonal inverse inductive matrix of the full system, the dependence of the coupler node fluxes on the qubit current $\tilde{I}_{\mathrm{q_{A}}}$ can be determined. The voltage of the second qubit $\tilde{V}_{\mathrm{q_{B}}}$ is then expressed in terms of the coupler fluxes, allowing the construction of the impedance between the qubits, $Z_{\mathrm{q_{A}q_{B}}}=\tilde{V}_{\mathrm{q_{B}}}/\tilde{I}_{\mathrm{q_{A}}}$. Following the methodology outlined in [54, 55], the XX coupling is defined as a function of the impedance between qubits:
$\begin{split}J_{\mathrm{q_{A}q_{B}}}=-\frac{1}{4}\sqrt{\frac{\omega_{\mathrm{q_{A}}}\omega_{\mathrm{q_{B}}}}{L_{\mathrm{q_{A}}}L_{\mathrm{q_{B}}}}}\textnormal{Im}\left[\frac{Z_{\mathrm{q_{A}q_{B}}}(\omega_{\mathrm{q_{A}}})}{\omega_{\mathrm{q_{A}}}}+\frac{Z_{\mathrm{q_{A}q_{B}}}(\omega_{\mathrm{q_{B}}})}{\omega_{\mathrm{q_{B}}}}\right],\end{split}$ (1)
where $\omega_{\mathrm{q_{A}}}$ and $\omega_{\mathrm{q_{B}}}$ are qubit mode frequencies, $L_{\mathrm{q_{A}}}$ and $L_{\mathrm{q_{B}}}$ are the total inductances of 3-JJ qubit circuits in this case. The total XX coupling between qubits is estimated as $J_{XX}=J_{\mathrm{q_{A}q_{B}}}+J_{0}$, where $J_{0}$ represents the direct capacitive interaction between the qubits. The imaginary part of the impedance here is
$\begin{split}\mathrm{Im}Z_{\mathrm{q_{A}q_{B}}}(\omega)=\omega\left(\frac{C_{\textnormal{23}}}{C_{\textnormal{23}}+C_{\mathrm{q_{A}}}}\right)^{2}\times\\ \times\frac{L_{\textnormal{J56}}\omega^{2}_{\mathrm{c_{+}}}\omega^{2}_{\mathrm{c_{0}}}\omega^{2}_{\mathrm{c_{-}}}}{(\omega^{2}_{\mathrm{c_{+}}}-\omega^{2})(\omega^{2}_{\mathrm{c_{0}}}-\omega^{2})(\omega^{2}_{\mathrm{c_{-}}}-\omega^{2})},\end{split}$ (2)
where $L_{\mathrm{J}56}$ is the flux-dependent effective inductance of the coupler SQUID, and $\omega_{c_{+}},\omega_{c_{0}},\omega_{c_{-}}$ are coupler normal modes frequencies depicted in Fig. 1(d). The mode indices indicate the relative directions of charge oscillations at the central coupler node with respect to the side nodes. Further details can be found in the Appendix B.
Once the XX coupling is known (see the green dashed-dotted line in Fig. 2(d)), the ZZ interaction strength can be determined. The Hamiltonian of the two-qubit system is constructed in the model of Duffing oscillators:
$\begin{split}H&=\omega_{\mathrm{q_{A}}}a^{\dagger}a+\frac{\delta_{\mathrm{q_{A}}}}{2}a^{\dagger}a(a^{\dagger}a-1)+\omega_{\mathrm{q_{B}}}b^{\dagger}b+\\ &+\frac{\delta_{\mathrm{q_{B}}}}{2}b^{\dagger}b(b^{\dagger}b-1)+J_{XX}(a^{\dagger}b+ab^{\dagger}).\end{split}$ (3)
The numerical Hamiltonian is constructed using the energy levels $|fg\rangle$, $|gf\rangle$, and $|ee\rangle$, while accounting for the XX interaction between levels. The ZZ interaction strength is determined as the frequency difference between the $|ee\rangle$ energy level in the undressed and dressed Hamiltonians. The results of this calculation are compared with experimental data and numerical approach in Fig. 2(d). Additional details are provided in Appendix B. A comparison between the proposed coupler and the single-mode tunable coupler is presented in Appendix C. An analysis of the scheme involving highly detuned qubits mediated by the three-mode coupler is provided in Appendix D.
III.3 CZ gate concept